# Generating crossmodal gene expression from cancer histopathology improves multimodal AI predictions

### v3

@@ 2.4.2 Visualisation of how transcriptomic data map to histopathology images
To add explainability, we compare the co-attention maps between transcriptomic data and the corresponding WSI patches, obtained using real and synthesized gene expression levels. TCGA-GBMLGG obtains a Spearman correlation of 0.939 and MAE of 2.07 and TCGA-KIRC obtains a Speraman correlation of 0.851 and MAE of 0.59 (Figures 2.a.iv, 2.b.iv and supplementary Table 3). A higher correlation with a high corresponding MAE would denote that the range of co-attention values is large for the particular gene group but the co-attention map is still correlated. For both cohorts, co-attention maps attributing to oncogenes are most correlated followed by tumour suppressor genes and transcription factor genes (Figures 2.a.iv, 2.b.iv and supplementary Table 3). Oncogenes usually drive cell division and survival, thus promoting cancer, inactivation of tumour suppressor genes can lead to cancer by failing to regulate uncontrolled cell growth Hanahan and Weinberg [2000], so the general co-attention observations above are not unexpected, although a rigorous inference would need a detailed analysis of gene set functional significance, which is outside the scope of this work.
@@
Further, we study the contribution of synthesized transcriptomic features in co-attending to the WSI patch embeddings. We observe that there are variances in the percentage contribution of individual gene groups in co-attention across both cohorts (Figures 2.a.v,2.b.v). For both cohorts, the transcription factor genes and oncogenes have the highest co-attention values with the WSI patches for both real and synthesized gene expression levels and hence contribute the most to the prediction (Figures 3.d, 2.a.v, 4.d, 2.b.v and supplementary Figure 10.d, 11.d). Activation of oncogenes is known to drive tumour proliferation Shay [1999], hence having these as one of the higher contributors to co-attention would be aligned with clinical expectations.
@@
We next visualise specific regions of the WSI and compare the distributed grades, distributed risk and the co-attention maps between synthesized transcriptomic features and the WSI patches (Figures 3.e, 4.e and supplementary Figure 10.e, 11.e). We observe that ground truth grade regions correspond to intermediate to high attention scores of transcription factors, and lower-grade regions are highly related to low attention values, as evident from the regions highlighted by the blue bounding box (Figures 3.c, 4.c).
Figure 5: Uncertainty quantification on the TCGA-GBMLGG dataset. (a) Performance of survival risk estimation and gradation for various demographic groups using real and synthesized transcriptomic data. (i) comparison between prediction using real and synthesized data (ii) performance for age groups (iii) performance for gender categories (iv) performance for survival status (censorship) categories (b) Plot of overall prediction results for fairness evaluation.
2.5 Uncertainty quantification for clinical reliability
2.5.1 Quantifying uncertainty
@@
We observe that the uncertainty profile of using synthesized transcriptomic data matches with that of using real transcriptomic data for gradation and survival risk estimation, for different patient demographics and stratification across both cohorts, by performing conformal prediction (CP) as described in Methods. For the entire test set population in TCGA-GBMLGG cohort, gradation uncertainty increases from 0.407 to 0.413 when synthesized data is used and the survival risk estimation uncertainty reduces from 0.446 to 0.438 (Figure 5.a.i). For TCGA-KIRC, gradation uncertainty increases from 0.292 to 0.334 and the survival risk uncertainty increases from 0.898 to 0.909 (Figure 6.a.i). However, the changes in uncertainty are not significant proving the similarity in uncertainty profile (Wilcoxon rank-sum test $p$-value $>0.05$, supplementary Tables 4, 5). Similarly for the different demographic groups for both cohorts, the uncertainty profile remains similar for real and synthesized transcriptomic data with no significant difference in uncertainty (Wilcoxon rank-sum test $p$-value $>0.05$, supplementary Tables 4, 5). It is observed that for a few population groups across both cohorts, the required coverage has not been met, because of fewer samples in the calibration set. But for many such cases, using synthesized data has increased the coverage at the cost of an insignificant increase in uncertainty (Wilcoxon rank-sum test $p$-value $>0.05$, Figures 6.a.i, 6.a.ii). As risk estimation directly uses the transcriptomic data for prediction and not just for co-attention, such differences in obtained coverage are expected even for minimal differences in the gene expression levels between the synthesized and real transcriptomic features. Thus, the uncertainty estimates help to understand the extent to which we may rely on the predictions, like the survival risk bounds. Additionally, the cancer grade conformal sets provide a range of alternative grades to consider before making the final decision.
Figure 6: Uncertainty quantification on the TCGA-KIRC dataset. (a) Performance of survival risk estimation and gradation for various demographic groups using real and synthesized transcriptomic data. (i) comparison between prediction using real and synthesized data (ii) performance for age groups (iii) performance for gender categories (iv) performance for survival status (censorship) categories (b) Plot of overall prediction results for fairness evaluation.
2.5.2 Fairness of predictions
@@
Next, we investigate algorithmic fairness, by assessing whether the uncertainty (normalised prediction set size) in gradation and survival risk predictions shows demographic biases. We observe that the predicted risk is less for patients who are reported to be alive and more for deceased patients, which is as expected (Figures 5.b, 6.b). Also, the uncertainty of survival risk estimation decreases with increasing risk, which is expected because the actual survival time of patients who are alive is beyond the reported survival time. Thus the model estimated survival risk should correspond to a time beyond the reported survival time and have a higher uncertainty measure when compared in terms of reported survival time. It may also be observed that the model tends to correlate higher risks with the highest grade, grade IV, for both cohorts. However, the trend of gradation uncertainty is different for the two cohorts. For TCGA-GBMLGG, the uncertainty of grade prediction is more for grade III, least for grade IV, and intermediate for grade II (supplementary Table 4 and Figure 5.b). But for TCGA-KIRC, the uncertainty is less for grade I and grade II with low estimated survival risks but high for all other grades (supplementary Table 4 and Figure 6.b). In general, the uncertainty of survival risk estimation is always greater if the survival time reported for a living patient is less (Figures 5.b, 6.b). Also, the uncertainty of prediction for alive patients is always high, except for gradation for TCGA-KIRC (Figures 5.a.iv, 6.a.iv). The uncertainty of grade prediction is more for male patients than females for both cohorts, whereas, for survival risk estimation, the uncertainty is more for females in TCGA-GBMLGG and males in TCGA-KIRC (Figures 5.a.iii, 6.a.iii). The uncertainty of risk estimation decreases with age for both cohorts (Figures 5.a.ii, 6.a.ii). For TCGA-GBMLGG, the uncertainty of grade prediction increases with decreasing magnification but remains similar for all magnification levels for TCGA-KIRC where magnification level 1 corresponds to the highest magnification and level 3 the lowest (supplementary Table 4). Thus these results provide an analysis from a health equity and predictive fairness point of view as to how the performance varies across patient demographics.
3 Methods
@@
The paper aims to use synthesized transcriptomic data for the automated selection of whole slide image (WSI) features for risk estimation and gradation of cancer cases. The proposed methodology is illustrated in Figure 1 of the main paper. The WSI is patchified and the corresponding patch embedding is obtained using UNI Chen et al. [2024]. Our novel model, PathGen, generates transcriptomic features from the patch embeddings. Together the patch embeddings and the gene embeddings obtained from the transcriptomic features are thus used by MCAT_GR for gradation and risk estimation. In the following subsections, we discuss the steps of the proposed methodology in detail.
3.1 Preprocessing
@@
Histopathology data. The diagnostic WSIs for a patient are divided into 224 $\times$ 224 patches, as required for obtaining patch embeddings using the pre-trained foundation model, UNI. The magnification level of the WSI is randomly chosen from the available options to make the model magnification level agnostic. The patches whose mean intensity normalized between 0 and 1 is more than 0.8 are excluded. It is observed that such patches consist mostly of the background region and hence, they do not contribute to the prediction. The 224 $\times$ 224 patches are provided as input to UNI to obtain the patch embeddings of dimension 1 $\times$ 1024. UNI Chen et al. [2024] is a pre-trained vision encoder for histopathology developed using private datasets, and thus is not based on the public datasets used in our experiments.
@@
Transcriptomic data. As proposed in MCAT Chen et al. [2021], the corresponding transcriptomic data of the patient is divided into 6 broad categories of genes - tumour suppressor genes, oncogenes, protein kinases, cell differentiation markers, transcription factors, and cytokines and growth factors. The gene expression levels are z-score normalized if already not done. Our model, PathGen, is trained to synthesize such transcriptomic data from the corresponding WSI. Embeddings of dimension 1 $\times$ 1024 are obtained for the gene expression levels of each category using the gene encoders.
Figure 7: Model Architecture. (a) Architecture of our novel diffusion model, PathGen, used for synthesizing gene expression levels from histopathology images. (b) Architecture of the modified MCAT model for gradation and risk estimation, MCAT_GR. (c) Detailed architecture of PathGen transformer. (d) Architecture of gene encoder that converts gene expression level to gene embeddings. (e) Representation of genomic-guided co-attention function. (f) Architecture of gene decoder that converts gene embeddings back to gene expression levels.
3.2 Synthesizing transcriptomic data from histopathology images
@@
The process of diffusion is represented as a Markov chain where in the forward process noise is added in each step to degrade the the data for a predetermined timestep $T$ Sohl-Dickstein et al. [2015]. Diffusion models are trained to predict the noise introduced in each step from a noisy sample. In the reverse process, the noise is given as input to the diffusion model and denoising is performed for $T$ timesteps to obtain the synthesized data Ho et al. [2020].
@@ 3.2.1 Architecture of PathGen
The architecture of our model, PathGen, to generate transcriptomic features from whole slide images is illustrated in Figure 7.a with further details in Figures 7.c, 7.d, 7.e and 7.f. To obtain the transcriptomic features at a timestep $t-1$, $\mathbf{x}_{t-1}$, the transcriptomic features at timestep $t$ and the WSI path embeddings are provided as input. The process is performed for T timesteps, where $\mathbf{x}_{T}$ is noise and $\mathbf{x}_{0}$ is the synthesized transcriptomic data. The embedding of the gene expression levels in the transcriptomic data at $\mathbf{x}_{t}$ is obtained using the gene encoder consisting of four linear layers with ELU activation as illustrated in Figure 7.d. The gene embedding and the corresponding patch embedding obtained using UNI Chen et al. [2024] goes as input to the PathGen transformer, illustrated in Figure 7.c. The PathGen transformer comprises genomic-guided co-attention illustrated in Figure 7.e., proposed in MCAT Chen et al. [2021], and the transformer encoder layers Vaswani et al. [2017]. Genomic-guided co-attention is inspired by the self-attention introduced in transformer Vaswani et al. [2017]. The gene embeddings attend to the WSI embeddings to produce co-attended embeddings which comprises of the most relevant combined features learned while training. Co-attention is performed thrice to ensure that the generated transcriptomic features have correspondence to the input WSI. Further, the transformer embeddings are decoded to gene expression levels using specific gene decoders for the different gene groups. The gene decoders comprise linear layers to ELU activation except the last, as illustrated in Figure 7.f.
3.2.2 Training and inference
@@
Algorithm 1 gives the training algorithm. The model is trained with the loss function in equation 7 for $T=1000$ timesteps with a learning rate of $1\times 10^{-4}$. For training a sample is considered at a random intermediate timestep $t$ sampled uniformly. Algorithm 2 specifies the inference algorithm. The inference is performed for $T=1000$ timesteps by initially providing random noise sampled from mean zero, unit variance Gaussian distribution as input using equation 4.
Algorithm 1 PathGen Training
repeat
     $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$
     $t\sim\text{Uniform}(\{1,\dots,T\})$
     $\epsilon\sim\mathcal{N}(0,\mathbf{I})$
@@
     Take gradient descent step on $\nabla_{\theta}\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}$
until converged
Algorithm 2 PathGen Inference
$\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$
for $t=T,\dots,1$ do
     $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$ if $t>1$, else $\mathbf{z}=0$
@@
     $\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z}$
end for
return $\mathbf{x}_{0}$
3.3 Gradation and risk estimation
@@
The synthesized transcriptomic data using our model, PathGen, along with the corresponding WSIs is used for gradation and risk estimation, using the model MCAT_GR.
3.3.1 Architecture
@@
The architecture used for gradation and risk estimation using multi-modal histopathology and synthesized transcriptomic data is MCAT Chen et al. [2021] with an additional pipeline for gradation. The modified MCAT architecture, MCAT_GR, is illustrated in Figure 7.b. The co-attended features processed using the pathomic transformer of MCAT_GR are used for gradation by passing through a separate global attention pooling. However, transcriptomic feature embeddings from the gene transformer are not used directly for gradation as we experiment to see that not including the transcriptomic data for grade prediction performs better. The global attention pooling performed for risk estimation and gradation are separate to decrease the dependency between the grade and risk.
3.3.2 Training details
@@
The loss used for training gradation is binary cross entropy loss Mao et al. [2023], $L_{grade}$, and the loss used for training risk estimation is negative log-likelihood survival loss Chen et al. [2021], $L_{risk}$. Thus, the total loss for MCAT_GR is given by the equation,
@@
where $\lambda$ is a hyperparameter to control the contribution of gradation and survival loss in the training. The learning rate used is $2\times 10^{-4}$.
3.4 Conformal Prediction and uncertainty estimation
@@
Conformal prediction is performed for gradation to provide the clinicians with a calibration set of the most probable grade options. Marginal conformal prediction ensures that the coverage of prediction obtained for a population is more than $1-\alpha$, where $\alpha$ is a chosen error rate Angelopoulos and Bates [2021]. The conformal condition for gradation is given by the equation,
@@
where $C$ represents the conformal set consisting of probable prediction classes to gain the desired coverage of at least $1-\alpha$. Algorithm 3 specifies the procedure for performing conformal gradation Angelopoulos and Bates [2021]. The uncertainty for gradation is computed using the conformal set by the equation,

### v4

@@ RESULTS
To add explainability, we next study how transcriptomic data co-attends to the corresponding WSI patched. We compare the co-attention maps obtained using real and synthesized gene expression levels using Spearman correlation coefficient and normalized mean absolute error (nMAE). A higher correlation with a high corresponding nMAE would denote that the range of co-attention values is large for the particular gene group but the co-attention map is still correlated. TCGA-GBMLGG obtains a correlation of 0.939 and nMAE of 0.043, TCGA-KIRC obtains a correlation of 0.851 and nMAE of 0.063, TCGA-UCEC obtains a correlation of 0.997 and nMAE of 0.022, TCGA-BRCA obtains a correlation of 0.992 and nMAE of 0.014, CPTAC-GBM obtains a correlation of 0.993 and nMAE of 0.013, and CPTAC-UCEC obtains a correlation of 0.998 and nMAE of 0.021 over all gene groups (Figure 5.b.i). These significantly high correlation values (Spearman correlation $p$-value <0.05<0.05) and low nMAEs between co-attention maps for real and synthesized transcriptomic data explains the reason for no significant difference between using the same for gradation and survival risk estimation across cohorts and datasets (Wilcoxon ranksum test $p$-value >0.05>0.05, ANOSIM test $p$-value >0.05>0.05, Supplementary Table LABEL:s-result_main). For all cohorts, co-attention maps attributing to oncogenes are most correlated with a low nMAE, followed by transcription factor genes and tumour suppressor genes (Figures 5.b.ii, 5.b.iii). Oncogenes usually drive cell division and survival, thus promoting cancer. Inactivation of tumour suppressor genes can lead to cancer by failing to regulate uncontrolled cell growth Hanahan and Weinberg [2000]. So the general co-attention observations above are not unexpected, although a rigorous inference would need a detailed analysis of gene set functional significance, which is outside the scope of this work.
@@
We further study the contribution of synthesized transcriptomic features in co-attending to the WSI patch embeddings and observe that there are variances in the percentage contribution of individual gene groups in co-attention for all cohorts and datasets (Figures 5.c.i,5.c.ii, and Supplementary Table LABEL:s-result_coattn_perc). For most cohorts, the transcription factor genes and oncogenes have the highest co-attention values with the WSI patches for both real and synthesized gene expression levels and hence contribute the most to the prediction (Figures 3.d, 4.d, and Supplementary Figures LABEL:s-fig_gbmlgg1.d, LABEL:s-fig_kirc1.d, LABEL:s-fig_cptac_ucec.d). Since activation of oncogenes is known to drive tumour proliferation Shay [1999], having these as one of the higher contributors to co-attention would be aligned with clinical expectations.
@@
We next visualise specific regions of the WSI and compare the distributed grades, distributed risk and the co-attention maps between synthesized transcriptomic features and the WSI patches (Figures 3.e, 4.e and Supplementary Figure LABEL:s-fig_gbmlgg1.e, LABEL:s-fig_kirc1.e, LABEL:s-fig_cptac_ucec.e). We observe that ground truth grade regions correspond to higher attention scores and other grade regions correspond to low attention values (Figures 3, 4, 5.a.ii).
Uncertainty quantification for clinical reliability
[…]
@@
It may be observed that for a few population groups across some cohorts, the obtained coverage does not always match the desired coverage because of fewer samples in the calibration set. These coverage slacks are expected and can be theoretically computed as a function of the cardinality of the calibration set Angelopoulos and Bates [2021]. We find that all obtained coverages that are less than the desired coverage of 0.9 fall within the expected deviation for the corresponding calibration set size of the demographic group (Supplementary Tables LABEL:s-data_demo, LABEL:s-table_deviation). For many such cases, using synthesized data has increased the coverage at the cost of an insignificant increase in uncertainty (Wilcoxon rank-sum test $p$-value >0.05>0.05, Supplementary Tables LABEL:s-result_GBMLGG, LABEL:s-result_GBM, LABEL:s-result_KIRC, LABEL:s-result_TUCEC, LABEL:s-result_CUCEC). As risk estimation directly uses the transcriptomic data for prediction and not just for co-attention, such differences in obtained coverage are expected even for minimal differences in the gene expression levels between the synthesized and real transcriptomic features. Moreover, the uncertainty estimates are derived from conformal sets for gradation and risk bounds for survival estimation, which helps to understand the extent to which we may rely on the model predictions. Additionally, the gradation conformal sets provide a range of alternative grades to consider before making the final decision.
@@
Next, we investigate algorithmic fairness by assessing whether the performance in gradation and survival risk predictions shows demographic biases. We observe that the predicted risk is less for patients who are reported to be alive and more for deceased patients, which is as expected (Figures 6.a, 6.c, 6.d). Also, the uncertainty of survival risk estimation decreases with increasing risk (Figures 6.a, 6.b, 6.c), which is expected because the actual survival time of patients who are alive is beyond the reported survival time. Thus, the model estimated survival risk should correspond to a time beyond the reported survival time and have a higher uncertainty measure when compared in terms of reported survival time. It may also be observed that the model tends to correlate higher risks with the highest grade, for all data cohorts (Figure 6). In general, the uncertainty of survival risk estimation is always greater if the survival time reported for a living patient is less (Figure 6). Also, the uncertainty of prediction for alive patients is mostly high, except for gradation in TCGA-KIRC (Figure 6 and Supplementary Tables LABEL:s-result_GBMLGG, LABEL:s-result_GBM, LABEL:s-result_KIRC, LABEL:s-result_TUCEC, LABEL:s-result_CUCEC). The uncertainty of grade and survival prediction is more for male patients than females except for gradation in TCGA-GBMLGG (Supplementary Tables LABEL:s-result_GBMLGG, LABEL:s-result_GBM, LABEL:s-result_KIRC). The uncertainty of risk estimation decreases with age for all TCGA cohorts (Supplementary Tables LABEL:s-result_GBMLGG, LABEL:s-result_KIRC, LABEL:s-result_TUCEC). Thus, these results provide an analysis from a health equity and predictive fairness point of view as to how the performance varies across patient demographics.
DISCUSSION
In this paper, we present PathGen, a crossmodal generative diffusion model for synthesizing transcriptomic features from WSIs of H&\&E stained slides. Our results are based on assessment of two publicly available cancer datasets - TCGA and CPTAC, using data from four cancer cohorts: glioma-glioblastoma, renal, uterine and breast. We demonstrated that PathGen synthesized transcriptomic features are highly correlated to true patient transcriptomes. By using a multimodal AI model to predict cancer grade and survival risk from WSIs and transcriptomic features, with state-of-the-art performance, we demonstrated that combining the features learnt from WSIs using a vision transformer based foundation model with the synthesized transcriptomic features from PathGen, significantly improved predictive accuracy compared to WSIs alone. Remarkably, we found that there was no significant difference in accuracy when real patient transcriptomic data were used instead of PathGen synthesized transcriptomic data in the multimodal model, demonstrating that PathGen captures the independent gradation/survival information in true patient gene expressions. To further check for generalizability, the models trained on the TCGA dataset were evaluated on the same CPTAC cohorts. We observed that our claims that synthesized transcriptomic data provides significant added value over WSI-only predictions, and that predictions based on real and synthesized transcriptomic data remain significantly close, generalizes well across independent datasets and cohorts.
H&\&E stained slides of tumour tissue are essential for diagnostic and prognostic assessment of malignancy, and are collected routinely. However, transcriptomic tests, though they have the potential to inform prognosis and guide treatment selection Sparano et al. [2018], Tsakiroglou et al. [2023] are not routinely performed, especially in the public sector Harnan et al. [2019]. This is due to a number of reasons, including economic cost, infrastructural requirements and incomplete knowledge about patient benefit. PathGen can synthesize realistic and informative patient transcriptomic features, from routine and inexpensive WSIs of H&\&E stained slides. Also, we note the average inference time for cross-modal synthesis of transcriptomic data from WSIs using PathGen is 21.15 seconds, and the average time for estimating survival and grade is 0.71 seconds, which varies over the magnification level and the size of the WSI, as expected. Thus, in the future, we envision that PathGen can be studied as a low cost screening tool, to guide the targeted collection of transcriptomic data modalities, by potentially selecting patients who would benefit from transcriptomic assessment for more accurate diagnosis/prognosis. This would of course need a dedicated follow up study which would include patient response to treatment when selected in this manner, that would entail a separate clinical validation project.
Moreover, to be useful in the clinic, AI models must be trusted Banerji et al. [2023]. Thus, to add explainability to the model predictions, we have undertaken measures with design choices guided by emerging regulatory requirements (e.g., EU AI Act Articles 13–15) Commission [2021] and in consultation with a practising histopathologist based at University College Hospitals, London. Firstly, by employing co-attention tools for model transparency, we visualize the role of the gene groups and WSI patches in predicting survival and grades, and by mapping distributed gradation and survival risk estimates back onto the WSI images, we provide a window into intra-tumour heterogeneity. These facilitate additional visual presentations to the clinician experts to choose whether or not to draw any further intuition from them. Secondly, we have employed conformal prediction for assessment of model reliability, providing not simply point estimates for gradation and survival risk, but ranges that contain true values with a specified certainty. This analysis has also allowed us to assess model fairness by comparing model reliability across patient demographics such as age and gender. Importantly, we found that there was no significant difference in model reliability or fairness when using PathGen synthesized transcriptomic features compared to real patient transcriptomes. For reproducibility of research and ethos of open science, we have made the code fully available online through our GitHub repository Dey et al. [2025].
However, though the PatheGen synthesized transcriptomic features corresponding to the six considered gene groups (both synthesized and real) contain important diagnostic/prognostic information, they do not capture the full diversity of the transcriptome. There is likely a limit to how much of the full transcriptome can be synthesized from only knowledge of WSIs. Also, evaluating the clinically important but less abundant synthesized gene expressions would require substantial domain expertise. Hence, we do not claim that PathGen replaces the need for real transcriptomic data, but rather it provides a potential way of estimating the benefit of collecting the full data across the population. The increasing use of techniques such as spatial transcriptomics Williams et al. [2022] will provide insight into this question, as well as WSI patch level resolution of transcriptomics, to guide development of synthesized transcriptomic data with a greater number of features in the near future.
METHODS
@@
The paper aims to use synthesized transcriptomic data for the automated selection of whole slide image (WSI) features for risk estimation and gradation of cancer cases, and evaluate the value added by the use of synthesized transcriptomic data over WSI-only predictions and the difference in using real transcriptomic data instead (Figure 1). The WSIs are patchified and the corresponding patch embeddings are obtained using UNI Chen et al. [2024]. Our model, PathGen, generates transcriptomic features from the patch embeddings. Together, the patch embeddings and the gene embeddings obtained from the transcriptomic features are thus used by MCAT_GR for gradation and risk estimation. In the following subsections, we discuss the steps of the proposed methodology in detail.
Preprocessing multimodal data
@@
Histopathology data. The diagnostic WSIs for a patient are divided into 224 $\times$ 224 patches, as required for obtaining patch embeddings using the pre-trained foundation model, UNI. The magnification level of the WSI is randomly chosen from the available options to make the model magnification level agnostic. The patches having normalized mean intensity more than 0.8 are excluded. It is observed that such patches consist mostly of the background region and hence, they do not contribute to the prediction. We do not use stain normalization in preprocessing the WSIs in line with recent works that find no significant effect of using stain normalization as a preprocessing step Voon et al. [2023]. The 224 $\times$ 224 patches are provided as input to UNI to obtain the patch embeddings of dimension 1 $\times$ 1024. UNI Chen et al. [2024] is a pre-trained vision encoder for histopathology developed using private datasets, and thus is not based on the public datasets used in our experiments.
@@
Transcriptomic data. As proposed in MCAT Chen et al. [2021], the corresponding transcriptomic data of the patient is divided into 6 broad categories of genes - tumour suppressor genes, oncogenes, protein kinases, cell differentiation markers, transcription factors, and cytokines and growth factors. The gene expression levels are z-score normalized if not already done. Our model, PathGen, is trained to synthesize such transcriptomic data from the corresponding WSIs. The gene encoders are trained to obtain embeddings of dimension 1 $\times$ 1024 for the gene expression levels of each category.
Synthesizing transcriptomic data from histopathology images
@@
The process of diffusion is represented as a Markov chain where, in the forward process, noise is added in each step to degrade the the data for a predetermined timestep $T$ Sohl-Dickstein et al. [2015]. Diffusion models are trained to predict the noise introduced in each step from a noisy sample. In the reverse process, the noise is given as input to the diffusion model and denoising is performed for $T$ timesteps to obtain the synthesized data Ho et al. [2020].
@@
The architecture of our model, PathGen Dey et al. [2025], to generate transcriptomic features from whole slide images is illustrated in Figure 7.a with further details in Figures 7.c, 7.d, 7.e and 7.f. To obtain the transcriptomic features at a timestep t1t-1, $\mathbf{x}_{t-1}$, the transcriptomic features at timestep $t$ and the WSI path embeddings are provided as input. The process is performed for T timesteps, where $\mathbf{x}_{T}$ is noise and $\mathbf{x}_{0}$ is the synthesized transcriptomic data. The embedding of the gene expression levels in the transcriptomic data at $\mathbf{x}_{t}$ is obtained using the gene encoder consisting of four linear layers with ELU activation as illustrated in Figure 7.d. The gene embedding and the corresponding patch embedding obtained using UNI Chen et al. [2024] goes as input to the PathGen transformer, illustrated in Figure 7.c. The PathGen transformer comprises genomic-guided co-attention illustrated in Figure 7.e., proposed in MCAT Chen et al. [2021], and the transformer encoder layers Vaswani et al. [2017]. Genomic-guided co-attention is inspired by the self-attention introduced in the transformer Vaswani et al. [2017]. The gene embeddings attend to the WSI embeddings to produce co-attended embeddings, which comprise the most relevant combined features learned while training. Co-attention is performed thrice to ensure that the generated transcriptomic features have correspondence to the input WSI. Further, the transformer embeddings are decoded to gene expression levels using specific gene decoders for the different gene groups. The gene decoders comprise linear layers to ELU activation except the last, as illustrated in Figure 7.f.
Training and inference
@@
Box Training and inference gives the training algorithm. The model is trained with the loss function in equation 7 for $T=1000$ timesteps with a learning rate of $1\times 10^{-4}$. For training, a sample is considered at a random intermediate timestep $t$ sampled uniformly. Box Training and inference specifies the inference algorithm. The inference is performed for $T=1000$ timesteps by initially providing random noise sampled from mean mean-zero, unit-variance Gaussian distribution as input using equation 4. All experiments were performed with an NVIDIA A100 GPU, including training and inference.
@@
Box 1 — PathGen training algorithm repeat   $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$   $t\sim\text{Uniform}(\{1,\dots,T\})$   $\epsilon\sim\mathcal{N}(0,\mathbf{I})$   Take gradient descent step on $\nabla_{\theta}\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}$ until converged
@@
Box 2 — PathGen training algorithm $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$ for $t=T,\dots,1$ do   $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$ if $t>1$, else $\mathbf{z}=0$   $\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z}$ end for return $\mathbf{x}_{0}$
Multimodal cancer gradation and survival risk estimation
@@
The PathGen synthesized transcriptomic data along with the corresponding WSIs, is used for gradation and risk estimation, using the predictive model MCAT_GR.
Architecture
@@
The architecture used for gradation and risk estimation using multi-modal histopathology and synthesized transcriptomic data is MCAT Chen et al. [2021] with an additional pipeline for gradation. The modified MCAT architecture, MCAT_GR, is illustrated in Figure 7.b. The co-attended features processed using the pathomic transformer of MCAT_GR are used for gradation by passing through a separate global attention pooling. However, transcriptomic feature embeddings from the gene transformer are not used directly for gradation, as we experiment to see that not including the transcriptomic data for grade prediction performs better. The global attention pooling performed for risk estimation and gradation is separate to decrease the dependency between the grade and risk.
Training details
@@
The loss used for training gradation is binary cross entropy loss Mao et al. [2023] ($L_{grade}$), and the loss used for training risk estimation is negative log-likelihood survival loss Chen et al. [2021] ($L_{risk}$). Thus, the total loss for MCAT_GR is given by the equation,
@@
where $\lambda$ is a hyperparameter to control the contribution of gradation and survival loss in the training. We use $\lambda=0.3$ for our experiments. However, on unavailability of gradation annotations, models are trained for survival only using $L_{risk}$ ($\lambda=0$). The learning rate used is $2\times 10^{-4}$. The models were trained using an NVIDIA A100 GPU.
[…]
Uncertainty estimation using conformal prediction
@@
Conformal prediction is performed to estimate the reliability of model predicted grades and to provide clinicians with a conformal set of the most probable grade options. Marginal conformal prediction ensures that the coverage of prediction obtained for a population is more than $1-\alpha$, where $\alpha$ is a chosen error rate Angelopoulos and Bates [2021]. The marginal coverage property for gradation is given by the equation,
@@
where $C$ represents the conformal set consisting of probable prediction classes to gain the desired coverage of at least $1-\alpha$. While marginal calibration ensures desired coverage for the entire population, it fails to ensure the fairness of coverage stratified over demographic groups. For instance, given a desired coverage of 90%, female patients may have obtained 100% coverage while male patients may have obtained 80% coverage, still satisfying the marginal coverage property. Thus, we further ensure coverage stratified over groups for all demographic categories. The stratified conditional coverage property for gradation is given by the equation,
$1-\alpha\leq\text{I\kern-1.49994ptP}{\,(Y_{test}\in C(X_{test})\;|\;X_{test}\in g_{i}}),\quad\forall\,g_{i}\in G$ (10)
where G represents the demographic category (gender, age, etc.) and $g_{i}$ represents the groups (male, female, age <40<40, age >60>60, etc) in the category. Box Training details specifies the procedure for performing conformal prediction for gradation Angelopoulos and Bates [2021]. The uncertainty for gradation is computed using the predicted conformal set by the equation,